(8) Solve the equation and find exact solutions. log2 (x + 1) = 1 - log2 (x - 1)
Answer #1

Soph Given that, log (x+1)= 1-log (3) we have to solve the given equation, ire dog, (x+1)=1-log, (x-1) & log₂ (x+1)+ log (x-1This is the required solution.